OpenStudy (anonymous):

sketch the level curve f(x,y)=C where f(x,y)=-e^x+y and C=2

OpenStudy (anonymous):

Write: $$f(x,y)=2\\-e^x+y=2\\y=e^x+2$$hence, on the plane \(z=2\), our graph looks like this: